OpenStudy (anonymous):

Create your own two-step equation with All Real Numbers as the solution. Show the steps to solving the equation to prove the solution All Real Numbers.

OpenStudy (anonymous):

2x - 13 = 5 2x - 13 + 13 = 5 + 13(add 13 to both sides to get 2x by itself) 2x = 18 (divide both sides by 2 to get x by itself) 2x/2 = 18/2 x = 9 check... 2x - 13 = 5 2(9) - 13 = 5 18 - 13 = 5 5 = 5 (correct)

OpenStudy (precal):